| You wish to test the following claim (Ha) at a significance level ofα = 0.10. You obtain the following two samples of data. Sample #1 77.6 95.1 82.6 85.8 89.2 85.5 95.1 841 101.7 98.3 100.6 101.2 87.6 85.2 78.9 72.2 94.2 103.7 87.6 64.2 100.76.7 76.7 100.1 85.5 90.2 64.2 101.2 Sample #2 76.1 84.5 87.3 83.3 96.1 94.9 86.4 79.7 87.5 81.7 86.2 82.9 81.2 102.3 75.8 68.3 89 83.9 83.3 89.5 91.2 87.1 89.8 69.3 72.6 91.2 97.7 93.4 916 97.795.5 73.1 88.5 81.4 64.7 78.5 91.9 96.1 90 100.1 97.2 80 76.2 76.7 70.2 83.5 90 92.7 82.6 106.6 98.3 99.6 77.2 87.6 91.3 What is the test statistic for this sample? (Report answer accurate to three decimal places.) test statistic 1,186 What is the p-valuc for this sample? For this calculation, use the degrees of freedom reported from the technology yo are using. (Report answer accurate to four decimal places.) p-value-10 05 las
Answer #1

Hypothesis

     H0: µ1= µ2

     Ha: µ1> µ2

Test statistics= t=1.186

Degree of freedom = n1 + n2 -2 =45 + 38 - 2 =81

P-value =0.1195

Conclusion:

The p-value of independent two sample t-test is greater than 0.1. Hence we accept null hypothesis at 0.1 level of significance. i.e. µ1= µ2